# Bridging Graph and State-Space Modeling for Intensive Care Unit Length of Stay Prediction

### v1

@@ 4.1 Results
As shown in Table 1 and Figure 3, S2G-Net achieves the best performance on all 6 evaluation metrics. It obtains the highest $R^{2}$ ($0.43\pm 0.01$) with statistically significant improvements in $R^{2}$, MSE, and Kappa ($p<0.05$). Compared to the best-performing baseline GraphGPS, S2G-Net improves $R^{2}$ by approximately 7.5% and reduces log-MAPE by 4.1%. Our model also delivers competitive results on MSLE and MAD, indicating that integrating spatio-temporal graph model with domain-specific priors may offer benefits for clinical prediction tasks.
@@
Figure 3: Multi-dimensional analysis of S2G-Net and baselines. (a) Training cost vs. $R^{2}$; (b) Parameter count vs. $R^{2}$; (c) $R^{2}$ vs. MSE; (d) Overall ranking by $R^{2}$.
@@
Time window analysis indicates that longer observation windows provide meaningful performance gains, especially between 6 and 24 hours (Figure 4a, d). Beyond 24 hours, marginal improvements indicate diminishing returns, potentially attributable to noise accumulation or reduced signal correlation in long-range dependencies.
@@
SHAP-based analysis revealed clinically meaningful predictors (Figure 5). The high contribution of GCS language and motor scores is consistent with previous studies, confirming that neurological status is the ICU prognosis strong predictor Bastos et al. (1993). ICD codes related to acute respiratory failure (J96.2) and acute kidney injury (N17.0) also match common ICU admission diagnoses and mortality risk factors Park and Faubel (2021); Havaldar et al. (2024). The effect of ICU identifiers like MICU and SICU reflects differences in patient severity and treatment protocols by specialty. Its excellent calibration performance (ECE = 0.003) provides reliable predictive results for real-time clinical decision support.
@@ 5 Conclusion
In this work, we introduced S2G-Net, a novel dual-path neural architecture for ICU LOS prediction. Our model uniquely combines SSMs to capture temporal patient data and GNNs to represent inter-patient relationships, leveraging this as a geometric prior. The results demonstrated that this hybrid approach consistently outperformed existing baselines on the large-scale MIMIC-IV dataset, proving to be a highly effective and computationally efficient solution for a critical task in healthcare. Importantly, this research validates the importance of a holistic approach that considers both an individual patient’s trajectory and relationships within the patient population. Future work will focus on extending this framework to other clinical tasks and real-time decision support systems to further improve patient care and optimize hospital resources.
@@ Appendix C Evaluation Metrics
Mean Squared Logarithmic Error (MSLE) Pedregosa et al. [2011].
@@
Mean Absolute Deviation (MAD) Willmott and Matsuura [2005].
@@
Linear Weighted Kappa (Kappa) Cohen [1968].
@@
Table 4: Resource Consumption Comparison of Baseline Models.
@@ Appendix E Hyperparameter Sensitivity
We performed 75 independent trials using Optuna’s tree-structured Parzen estimator (TPE) Akiba et al. [2019]. Figure 6a shows that validation $R^{2}$ increases steadily across 75 optimization trials, improving rapidly at first and tapering after approximately 12 trials. The performance histogram (Figure 6b) demonstrates a unimodal distribution, suggesting stable search behavior and the absence of degenerate hyperparameter regions.

### v2

@@ 4.1 Results
As shown in Table 1 and Figure 3, S2G-Net consistently outperforms both neural and non-neural baselines across all 6 evaluation metrics. It obtains the highest $R^{2}$ ($0.43\pm 0.01$) with statistically significant improvements in $R^{2}$, MSE, and Kappa. Compared to the best-performing baseline GraphGPS, S2G-Net improves $R^{2}$ by around 7.5% and reduces log-MAPE by 4.1%. It also delivers competitive results on MSLE and MAD, indicating that integrating spatio-temporal graph model with domain-specific priors may offer benefits for clinical prediction tasks.
@@
Figure 3: Multi-dimensional analysis of S2G-Net and baselines. (a) Training cost vs. $R^{2}$; (b) Parameter count vs. $R^{2}$; (c) $R^{2}$ vs. MSE; (d) Overall ranking by $R^{2}$. XGBoost excluded for non-comparable computational metrics.
@@
Time window analysis reveals that longer observation windows provide substantial performance gains, especially between 6 and 24 hours (Figure 4a, d). Beyond 24 hours, the benefits plateau, possibly attributed to noise accumulation or reduced signal correlation in long-range dependencies.
@@
SHAP-based analysis revealed clinically meaningful predictors (Figure 5). The high contribution of GCS language and motor scores is consistent with previous studies, confirming that neurological status is the ICU prognosis strong predictor [2]. ICD codes related to acute respiratory failure (J96.2) and acute kidney injury (N17.0) also match common ICU admission diagnoses and mortality risk factors [47, 25]. The effect of ICU identifiers like MICU and SICU reflects differences in patient severity and treatment protocols by specialty. Its excellent calibration performance (ECE = 0.003) provides reliable predictive results for real-time clinical decision support.
@@ 5 Conclusion
In this work, we introduced S2G-Net, a novel dual-path neural architecture for ICU LOS prediction. Our model uniquely combines SSMs to capture temporal patient data and GNNs to represent inter-patient relationships, leveraging this as a geometric prior. The results demonstrated that this hybrid approach consistently outperformed existing baselines on the large-scale MIMIC-IV dataset, proving to be a highly effective and computationally efficient solution for a critical task in healthcare. Importantly, this research validates the importance of a holistic approach that considers both an individual patient’s trajectory and relationships within the patient population. Future work will include external validation of S2G-Net on independent cohorts to assess generalizability and extending it to broader clinical tasks and real-time decision support to enhance patient care and resource efficiency.
@@ Appendix C Evaluation Metrics
Mean Squared Logarithmic Error (MSLE) [48].
@@
Mean Absolute Deviation (MAD) [63].
@@
Linear Weighted Kappa (Kappa) [8].
@@
Table 4: Resource Consumption Comparison of Baseline Models. XGBoost is excluded as its computational metrics are not comparable to neural models.
@@ Appendix E Hyperparameter Sensitivity
We performed 75 independent trials using Optuna’s tree-structured Parzen estimator (TPE) [1]. Figure 6a shows that validation $R^{2}$ increases steadily across 75 optimization trials, improving rapidly at first and tapering after approximately 12 trials. The performance histogram (Figure 6b) demonstrates a unimodal distribution, suggesting stable search behavior and the absence of degenerate hyperparameter regions.